OpenStudy (anonymous):

How much work is done if it takes 50N to move an object 10m?

OpenStudy (anonymous):

Multiply them. After all, W=Fd so W=50*10 W=500

OpenStudy (anonymous):

F = 50 N d = 10 m \[W = Fd\]so W = (50)(10) = 500 J Answer: W = 500 J

OpenStudy (anonymous):

W = F x W = 50 x 10 W = 500 joule